# Energy-Efficient Port Selection and Beamforming Design for Integrated Data and Energy Transfer Assisted by Fluid Antennas

### v1

@@ III-C Transmit beamforming optimization
[…]
@@
where $\boldsymbol{A}=\eta\boldsymbol{G}^{H}\hat{\boldsymbol{a}}\hat{\boldsymbol{a}}^{H}\boldsymbol{G}$ and $\boldsymbol{B}=\boldsymbol{H}^{H}\hat{\boldsymbol{a}}\hat{\boldsymbol{a}}^{H}\boldsymbol{H}$. According to [23], the optimal solution $\boldsymbol{W}^{*}$ of (P7) satisfies $\text{rank}(\boldsymbol{W}^{*})\leq 3$. Since $\boldsymbol{W}^{*}\neq\boldsymbol{0}$, thus we have $\text{rank}(\boldsymbol{W}^{*})=1$, which indicates that the rank constraint relaxation is tight and (P7) has the same optimal solution as (P6). By performing eigenvalue decomposition over $\boldsymbol{W}^{*}$, the optimal beamforming vector $\boldsymbol{w}^{*}$ is obtained. Note that when $\beta$ is fixed, the objective value of (P3) is non-decreasing by solving (P5) and (P7) alternately. Thus the proposed AO algorithm can converge to a sub-optimal solution. The overall proposed algorithm for solving (P3) is summarized in Algorithm 2. However, the optimized solution $\hat{\boldsymbol{a}}^{*}$ does not inherently possess binary properties. In order to recover the original port activation indicators, we choose the index of largest value of $\boldsymbol{a}_{i}^{*}$ as the selected port.
@@
0:  Initialize the WET efficiency $\beta^{(0)}=0$ and generate the port activation indicator $\hat{\boldsymbol{x}}^{(0)}$ and the beamforming vector $\boldsymbol{w}^{(0)}$ randomly. Set the maximal iteration number $M$ and the tolerance threshold $\epsilon$ ;
0:  For a given $\beta^{(m-1)}$
@@
7:  Update $\beta^{(m)}=\frac{E_{h}^{(m)}}{E_{t}^{(m)}}$ and set $m=m+1$.
8:  Until converges, i.e. $\frac{\beta^{(m)}-\beta^{(m-1)}}{\beta^{(m)}}\leq\epsilon$
@@ IV-A Problem formulation and Transformation
[…]
@@
[…]
@@
where $\eta_{EE}(s_{t},a_{t})$,$E_{h}(s_{t},a_{t})$ and $R(s_{t},a_{t})$ denote the instantaneous WET efficiency, harvested energy and achievable throughput by taking action $a_{t}$ at state $s_{t}$, respectively. $\beta$ is the discount factor. In (P9), the expected long term discount WET efficiency is maximized by searching for the optimal port selection policy $\pi^{*}$, while the constraints of expected long term discount throughput and harvest energy are guaranteed. However, (P9) is hard to be directly solved by DRL algorithm due to the expected long-term discount constraints. Therefore, we transform (P9) into multi-objective MDP problem and use zero-sum Markov Bandit game to solve it [31]. Maximizing the objective of (P9) is equivalent to obtain the maximal value of $\delta$ by satisfying
@@
The maximum $\delta$ can be found by exploiting bisection search, where a feasible policy $\pi$ can be found by satisfying the constraints (28a) to (IV-A). When the objective value $\delta$ of (P10) is optimal, the corresponding policy $\pi$ is also the optimal policy. In order to solve (P10), we formulate the problem as zero-sum Markov-Bandit game under a fixed objective value $\delta$, where the DRL agent solves a Markov decision problem, while its opponents solve a Bandit optimization problem. The zero-sum Markov-Bandit game is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{O}=\{o_{1},o_{2},o_{3}\}$ denote the opponents’ space, which corresponds to the expected long-term discount constraints of (P10). $\mathcal{P}$ represents the transition probability function among different states and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathcal{R}$ is the expected reward function. The zero-sum Markov-Bandit game is aimed to maximize the total discounted reward given by
@@ IV-B1 DRL definitions
$R(s_{t},a_{t},o)=\left\{\begin{aligned} &\eta_{EE}(\boldsymbol{s}_{t},\boldsymbol{a}_{t})-\overline{\delta},\qquad o=o_{1},\\ &E_{h}(\boldsymbol{s}_{t},\boldsymbol{a}_{t})-\tilde{E}_{th},\qquad o=o_{2},\\ &R(\boldsymbol{s}_{t},\boldsymbol{a}_{t})-\tilde{R}_{th},\qquad o=o_{3},\\ \end{aligned}\right.$ (32)
@@ IV-B2 Framework and updating process of the C-SAC algorithm
$\pi^{*}=\text{arg}\max_{\pi}\mathbf{E}_{(\boldsymbol{s}_{t},\boldsymbol{a}_{t})\sim\rho_{\pi}}\left[\sum_{t=1}^{\infty}\beta^{t-1}\left(R(\boldsymbol{s}_{t},\boldsymbol{a}_{t})\right.\right.$
@@
where $\alpha$ is the temperature parameter to control the stochasticity of the policy by adjusting the importance of the entropy term; $\mathcal{H}(\pi(\cdot|s_{t}))=\mathbf{E}_{a_{t}}\left[-\log\pi(a_{t}|s_{t})\right]$ is the policy entropy. Therefore, when the SAC algorithm is adopted to tackle the zero-sum Markov-Bandit game, the optimal policy $\bar{\pi}^{*}$ can be obtained via
$\bar{\pi}^{*}=\text{arg}\max_{\pi}\min_{o}\mathbf{E}_{(\boldsymbol{s}_{t},\boldsymbol{a}_{t})\sim\rho_{\pi}}\left[\sum_{t=1}^{\infty}\beta^{t-1}\left(R(\boldsymbol{s}_{t},\boldsymbol{a}_{t},o)\right.\right.$
@@
In order to obtain the optimal port selection policy, the SAC algorithm uses stochastic gradient descent (SGD) to update the deep neural network (DNN) at each training step. Two group of critic networks have the same DNN structure, and each group of critic network consists of a critic evaluate network and a critic target network. These critic networks can be denoted as $Q_{\theta_{1}}$, $Q_{\bar{\theta}_{1}}$, $Q_{\theta_{2}}$ and $Q_{\bar{\theta}_{2}}$, where $\theta_{n},\bar{\theta}_{n},n=\left\{1,2\right\}$ are the DNN weights of critic evaluate networks and critic target networks, respectively. The parameters of each critic network is updated by minimizing the soft Bellman residual, which can be represented as
@@
$J_{\pi}(\phi)=\mathbf{E}_{\boldsymbol{s}_{t}\sim\mathcal{D}}\left[\mathbf{E}_{\boldsymbol{a}_{t}\sim\pi_{\phi}}\left[\alpha\log\pi_{\phi}(\boldsymbol{a}_{t}|\boldsymbol{s}_{t})\right.\right.$
@@
Wavelength ($\lambda$) 6cm Noise variance $(\sigma^{2})$ $10^{-8}$
@@
Slack penalty $p$ 1000 Circuit power 0.05W
@@ V Simulation Results
Batch size($B_{s}$) 256 Target entropy $(\mathcal{H}_{0})$$N_{t}$
Replay buffer length ($L_{buffer}$) 20000 Average achievable data rate constraint $(\tilde{R}_{th})$ $8$bit/s/Hz
@@
In this section, we evaluate both the short term and long term WET efficiency of multiple FAs assisted IDET system by using simulation results. The distance from transmitter to DR and ER are set to be 10 meters and 3 meters, respectively. The signal attenuation at a reference distance of 1 meter is set as 30 dB. The path-loss factor of the both links are set to be 2.2. The whole model of our C-SAC algorithm is built on the platform of Pytorch 2.5.1 in Python 3.9. Both of the actor network and the critic networks have the dense layer 512$\times$256. The ReLU function is used for the hidden layers of both networks, while the Tanh and Linear functions are applied to the output layers of the actor and critic networks, respectively. The spatial and temporal correlation channel data is generated by Eq. (II-B) and Eq. (II-B). The temporal correlation factor is set as 0.98. Without specific statement, other default parameters about FAS assisted IDET system and and C-SAC algorithm are provided in Table I and in Table II. Note that those parameters are set for the sake of presentation. Using different parameters will lead to a shifted system performance, but with the same observations and conclusions.
@@ V-A Evaluation of short term performance
Fig. 3 depicts that the trade-off between WET efficiency and the transmission throughput requirement at DR with different number of FAs. Meanwhile, the WET efficiency is compared between our proposed scheme and a benchmark, which maximizes the harvested energy at ER by optimizing the beamforming vector and port activation indicator alternately. The voltage difference $\Delta\phi$ and moving power $P_{u}$ are set as 0.2 V and 2 W, respectively. Observe from Fig. 3 that the WET efficiency decreases with the transmission throughput requirement at DR.This is because a higher throughput requirement demands more transmit power for the DR, reducing the energy harvested by the ER. In addition, with the increment of transmitting FAs, the WET efficiency becomes higher. This is due to the additional antennas providing higher degrees of freedom, enabling more efficient wireless energy transfer. Moreover, the WET efficiency of our proposed scheme is much better than that of benchmark, which demonstrates that our proposed scheme can provide more efficient integrate data and energy transfer.
@@
Fig. 4 compares the WET efficiency of our scheme and the delay-insensitive scheme under different average velocity $u$ of FA. In the delay-insensitive scheme, port switching delay and energy consumption are not taken into account during the optimization process. Observe from Fig. 4 that the WET efficiency performance of our scheme outperforms that of delay-insensitive scheme. In the delay-insensitive scheme, FA movement is assumed to be instantaneous and free of any delay and energy consumption. However, in practice, this assumption results in excessively large movement steps, which in turn increase both delay and energy consumption, ultimately degrading the WET efficiency. In addition, with the increasing of the average velocity $u$ of FA, the WET efficiency of both schemes improves and the performance gap between two schemes is also narrowing, since a higher average velocity results in a lower port switching delay and energy consumption.
@@
Fig. 5 evaluates the average WET efficiency versus port number $N$ with different scaling constant $W$. The beamforming vector in traditional antenna assisted IDET system is optimized to maximize the WET efficiency under achievable throughput constraint and harvested energy constraint. The achievable throughput threshold is set as 7 bit/s/Hz. Observe from Fig. 5 that the average WET efficiency of FA assisted IDET system outperforms that of 4 traditional antenna counterparts and achieves performance comparable to that of 5 traditional antenna counterparts. This improvement is attributed to the higher degrees of freedom offered by FAs compared to traditional fixed antennas. In addition, the average WET efficiency increases with the port number, since a higher port number can catch more CSI of different position, which results in a better port selection. However, different from the previous analysis in [23] that a larger antenna size $W$ results in a better IDET performance, the average WET efficiency increases with $W$ initially and then decreases. This is because, although a larger antenna size reduces channel correlation, potentially leading to better channel quality, the delay and energy consumption will also increase accordingly. Therefore, selecting an appropriate antenna size is also crucial for improving the IDET performance when the switching delay and energy consumption of FA movement cannot be ignored.

### v2

@@ III-C Transmit beamforming optimization
$\text{(P7)}:\max_{\boldsymbol{W}}\quad$ $\text{tr}(\boldsymbol{AW})-q\text{tr}(\boldsymbol{W}),$ (24)
@@
where $\boldsymbol{A}=\eta\boldsymbol{G}^{H}\hat{\boldsymbol{a}}\hat{\boldsymbol{a}}^{H}\boldsymbol{G}$ and $\boldsymbol{B}=\boldsymbol{H}^{H}\hat{\boldsymbol{a}}\hat{\boldsymbol{a}}^{H}\boldsymbol{H}$. According to [26], the optimal solution $\boldsymbol{W}^{*}$ of (P7) satisfies $\text{rank}(\boldsymbol{W}^{*})\leq 3$. Since $\boldsymbol{W}^{*}\neq\boldsymbol{0}$, thus we have $\text{rank}(\boldsymbol{W}^{*})=1$, which indicates that the rank constraint relaxation is tight and (P7) has the same optimal solution as (P6). By performing eigenvalue decomposition over $\boldsymbol{W}^{*}$, the optimal beamforming vector $\boldsymbol{w}^{*}$ is obtained. Note that when $q$ is fixed, the objective value of (P3) is non-decreasing by solving (P5) and (P7) alternately. Thus the proposed AO algorithm can converge to the sub-optimal solutions. The overall proposed algorithm for solving (P3) is summarized in Algorithm 2. However, the optimized solution $\hat{\boldsymbol{a}}^{*}$ does not inherently possess binary properties. In order to recover the original port activation indicators, we choose the index of largest value of $\boldsymbol{a}_{i}^{*}$ as the selected port.
@@
0: Initialize the WET efficiency $q^{(0)}=0$ and generate the port activation indicator $\hat{\boldsymbol{x}}^{(0)}$ and the beamforming vector $\boldsymbol{w}^{(0)}$ randomly. Set the maximal iteration number $M$ and the tolerance threshold $\epsilon$ ;
0: For a given $q^{(m-1)}$
@@
7: Update $q^{(m)}=\frac{E_{h}^{(m)}}{E_{t}^{(m)}}$ and set $m=m+1$.
8:Until converges, i.e. $\frac{q^{(m)}-q^{(m-1)}}{q^{(m)}}\leq\epsilon$
@@
III-D Complexity analysis
First, we denote the iteration number of the FPP-SCA algorithm for solving (P5) as $I_{s}$, the iteration number of AO algorithm for solving (P3) as $I_{a}$ and the iteration number of Dinkelbach’s algorithm as $I_{d}$. According to [34] and [35], the computational complexity of FPP-SCA containing $x$ variables and $y$ constraints can be formulated as $\mathcal{O}(log(\frac{1}{\gamma})(x+y)^{3.5})$, while the computational complexity of SDP containing $x\times x$ positive semidefinite matrix and $y$ SDP constraints is $\mathcal{O}(\sqrt{x}log(\frac{1}{\gamma})(yx^{3}+x^{2}y^{2}+y^{3}))$, where $\gamma$ denotes the desired solution accuracy. Therefore, when the interior-point algorithm is applied to solve (P5) and (P7), the computational complexity of (P5) and (P7) can be expressed as $\mathcal{O}_{1}=\mathcal{O}(log(\frac{1}{\gamma})(NN_{t}+3)^{3.5})$ and $\mathcal{O}_{2}=\mathcal{O}(log(\frac{1}{\gamma})\sqrt{N_{t}}(3N_{t}^{3}+9N_{t}^{2}+27))$, respectively. The worst-case computational complexity of the AO algorithm for solving (P2) can be obtained as $\mathcal{O}(I_{d}I_{a}(I_{s}\mathcal{O}_{1}+\mathcal{O}_{2}))$.
@@ IV-A Problem formulation and Transformation
$\text{(P8)}:\max_{\boldsymbol{w}(t),\{k_{i}(t)\}}\quad$ $\tilde{\eta}_{EE}=\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\eta_{EE}(t),$ (25)
@@
$\text{(P9)}:$ $\max_{\pi}\quad\mathbf{E}_{\pi}\left[(1-\beta)\sum_{t=1}^{\infty}\beta^{t-1}\eta_{EE}(s_{t},a_{t})\right],$ (26)
@@
where $\eta_{EE}(s_{t},a_{t})$, $E_{h}(s_{t},a_{t})$ and $R(s_{t},a_{t})$ denote the instantaneous WET efficiency, harvested energy and achievable throughput by taking action $a_{t}$ at state $s_{t}$, respectively. $\beta$ is the discount factor. In (P9), the expected long term discount WET efficiency is maximized by searching for the optimal port selection policy $\pi^{*}$, while the constraints of expected long term discount throughput and harvested energy are guaranteed. However, (P9) is hard to be directly solved by DRL algorithm due to the expected long-term discount constraints. Therefore, we transform (P9) into multi-objective MDP problem and use zero-sum Markov Bandit game to solve it [37]. Maximizing the objective of (P9) is equivalent to obtain the maximal value of $\delta$ by satisfying
@@
In (P10), the discounted WET efficiency constraint prevents the agent from prioritizing short-term energy gains, which could lead to excessive port switching delay and future performance degradation. The discounted throughput constraint ensures that the average throughput stays above the required level, preventing the agent from overly favoring energy transfer at the cost of communication quality. These constraints influence the agent’s long-term decisions: a higher average energy target promotes the long-term WET efficiency, potentially sacrificing throughput, while a higher average throughput target prioritizes DR performance, which may reduce energy harvesting performance and WET efficiency of ER. The maximum $\delta$ can be found by exploiting bisection search, where a feasible policy $\pi$ can be found by satisfying the constraints (28a) to (IV-A). When the objective value $\delta$ of (P10) is optimal, the corresponding policy $\pi$ is also the optimal policy. In order to solve (P10), we formulate the problem as zero-sum Markov-Bandit game under a fixed objective value $\delta$, where the DRL agent solves a Markov decision problem, while its opponents solve a Bandit optimization problem. The zero-sum Markov-Bandit game is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{O}=\{o_{1},o_{2},o_{3}\}$ denotes the opponents’ space, which corresponds to the expected long-term discount constraints of (P10). $\mathcal{P}$ represents the transition probability function among different states and $\mathcal{R}:\mathcal{S}\times\mathcal{A}\to\mathcal{R}$ is the expected reward function. The zero-sum Markov-Bandit game is aimed to maximize the total discounted reward given by
@@ IV-B1 DRL definitions
[…]
@@ IV-B2 Framework and updating process of the C-SAC algorithm
[…]
@@
where $\alpha$ is temperature parameter to control the stochasticity of the policy by adjusting the importance of the entropy term; $\mathcal{H}(\pi(\cdot|s_{t}))=\mathbf{E}_{a_{t}}\left[-\log\pi(a_{t}|s_{t})\right]$ is the policy entropy. Therefore, when the SAC algorithm is adopted to tackle the zero-sum Markov-Bandit game, the optimal policy $\bar{\pi}^{*}$ can be obtained via
[…]
@@
In order to obtain the optimal port selection policy, the SAC algorithm uses stochastic gradient descent (SGD) to update the deep neural network (DNN) at each training step. Two group of critic networks have the same DNN structure, and each group of critic network consists of a critic evaluate network and a critic target network. These critic networks can be denoted as $Q_{\theta_{1}}$, $Q_{\bar{\theta}_{1}}$, $Q_{\theta_{2}}$ and $Q_{\bar{\theta}_{2}}$, where $\theta_{n},\bar{\theta}_{n},n=\left\{1,2\right\}$ are the DNN weights of critic evaluate networks and critic target networks, respectively. The parameters of each critic network are updated by minimizing the soft Bellman residual, which can be represented as
@@
[…]
@@
IV-C Complexity analysis
The computational complexity of the C-SAC algorithm primarily arises from two aspects: the training complexity and the execution complexity [41]. During the training phase, the gradient descent complexity for both the actor network and the critic network can be expressed as $\mathcal{O}\left(B_{s}\left(\sum_{k=0}^{K-1}n_{k}^{\pi}n_{k+1}^{\pi}+\sum_{j=0}^{J-1}n_{j}^{Q}n_{j+1}^{Q}\right)\right)$, where $B_{s}$ is the batch size, $K$ and $J$ represent the number of fully connected layers in the actor and critic networks, respectively. $n_{k}^{\pi}$ and $n_{j}^{Q}$ denote the number of neurons in the $k$-th layer of the actor network and the $j$-th layer of the critic network, respectively, with $k=0$ and $j=0$ referring to the input layers. Therefore, the overall complexity for training the C-SAC algorithm is $\mathcal{O}\left(T_{max}B_{s}\left(\sum_{k=0}^{K-1}n_{k}^{\pi}n_{k+1}^{\pi}+\sum_{j=0}^{J-1}n_{j}^{Q}n_{j+1}^{Q}\right)\right)$, where $T_{max}$ is the maximum number of training epochs. During the execution phase, only the trained actor network is required, and thus the computational complexity at each transmission frame corresponds to a forward pass through the actor network, which can be expressed as $\mathcal{O}\left(\sum_{k=0}^{K-1}n_{k}^{\pi}n_{k+1}^{\pi}\right)$.
@@ IV-C Complexity analysis
Wavelength ($\lambda$) 6cm Noise variance $(\sigma^{2})$ $10^{-8}$W
@@
Slack penalty pp 1000 Circuit power 0.05W
@@ V Simulation Results
Batch size($B_{s}$) 256 Target entropy $(\mathcal{H}_{0})$ -$N_{t}$
Replay buffer length ($L_{buffer}$) 20000 Average achievable data rate constraint $(\tilde{R}_{th})$ 88bit/s/Hz
@@
[…]
@@ V-A Evaluation of short term performance
To evaluate the effectiveness of our proposed scheme, three benchmark schemes are considered in this section:
  • Energy harvesting oriented scheme: The harvested energy at the ER is maximized by jointly optimizing the beamforming vector and port selection, subject to the achievable throughput constraint. The optimization problem can be solved by using our proposed AO algorithm. This scheme reflects an energy-centric design that prioritizes energy transfer without explicitly considering overall WET efficiency.
  • Delay-insensitive scheme: The WET efficiency is maximized by jointly optimizing the beamforming vector and port selection under the idealized assumption that port switching delay and power consumption are negligible, consistent with the common assumption in existing studies [26, 42]. Note that while these costs are ignored during optimization, they are still accounted for during performance evaluation.
  • Fixed-position antenna (FPA) scheme: In this scheme, the fluid antennas at the transmitter are replaced by conventional fixed-position antennas (FPA). Thus, port switching delay and power consumption are irrelevant. The WET efficiency is maximized by optimizing the beamforming vector under both throughput and harvested energy constraints. The Dinkelbach’s algorithm and SDR approach are adopted to obtain the optimal beamforming vector.
@@
Fig. 3 depicts that the trade-off between WET efficiency and the transmission throughput requirement at DR with different number of FAs. Meanwhile, the WET efficiency is compared between our proposed scheme and the energy harvesting oriented scheme. The voltage difference $\Delta\phi$ and moving power $P_{u}$ are set as 0.2 V and 2 W, respectively. Observe from Fig. 3 that the WET efficiency decreases with the transmission throughput requirement at DR. This is because under a higher throughput requirement, the optimization process prioritizes improving the signal quality at the DR to satisfy the stricter data rate constraint. As a result, the beamforming vector and port selection are steered more towards the DR, which may cause the transmission beam to become less favorable for the ER, thereby reducing the effective harvested energy. In addition, with the increment of transmitting FAs, the WET efficiency becomes higher. This is due to the additional antennas providing higher degrees of freedom, enabling more efficient wireless energy transfer. Moreover, the WET efficiency of our proposed scheme is much better than that of benchmark, which demonstrates that our proposed scheme can provide more efficient integrate data and energy transfer.
@@
Fig. 4 compares the WET efficiency of our scheme and the delay-insensitive scheme under different average velocity $u$ of FA. Observe from Fig. 4 that the WET efficiency of our proposed scheme consistently outperforms that of the delay-insensitive scheme. This is because that the delay-insensitive scheme assumes that the FA movement is instantaneous and incurs no switching delay or energy consumption during the optimization process, but the actual system evaluation still accounts for these physical effects. Since the delay-insensitive scheme ignores the cost of FA movement in decision-making, it tends to generate aggressive port switching with large movement distances. This results in significant delays and energy consumption in practice, ultimately degrading the real WET efficiency. In contrast, our proposed scheme explicitly incorporates the port switching delay and energy consumption into the optimization, effectively reducing unnecessary movement and achieving higher practical WET efficiency. Furthermore, as the average droplet velocity increases, both the port switching delay and movement energy consumption decrease for both schemes, leading to improved WET efficiency. The performance gap between the two schemes narrows at higher velocities because the switching cost becomes less significant. However, even at high velocities, our proposed scheme consistently outperforms the delay-insensitive scheme by achieving more energy-aware port selection. The performance gain highlights the novelty of our work, as we are the first to explicitly incorporate the port switching delay and energy consumption of FAs into the optimization process.
@@
Fig. 5 evaluates the average WET efficiency versus port number $N$ with different scaling constant $W$. In addition, the average WET efficiency is compared with that of FPA scheme. The achievable throughput threshold is set as 7 bit/s/Hz. As is shown in Fig. 5, when the FA has a port number $N=100$ and an antenna size $W=0.5$, the average WET efficiency of the FAS-assisted IDET system is 25% higher than that of its 4-antenna FPA counterpart and 4.3% higher than that of 5-antenna FPA system. This demonstrates that the FAS can achieve higher energy transfer efficiency with fewer antennas in IDET systems compared to conventional FPAs. This improvement is attributed to the additional spatial degrees of freedom provided by the FAS, which flexibly adjusts antenna positions to optimize energy transmission. In addition, the average WET efficiency increases with the port number, since a higher port number can capture more CSI from different positions, which results in a better port selection. However, different from the previous analysis in [26] that a larger antenna size $W$ results in a better IDET performance, the average WET efficiency increases with $W$ initially and then decreases. This is because, although a larger antenna size reduces channel correlation, potentially leading to better channel quality, the delay and energy consumption will also increase accordingly. Therefore, selecting an appropriate antenna size is also crucial for improving the IDET performance when the switching delay and energy consumption of FA movement cannot be ignored.